# On the nature of buckling instability in galactic bars

### v1

@@
Many strong simulated galactic bars experience buckling instability, which manifests itself as a vertical distortion out of the disk plane which later dissipates. Using a simulation of an isolated Milky Way-like galaxy I demonstrate that the phenomenon can be divided into two distinct phases. In the first one, the distortion grows and its pattern speed remains equal to the pattern speed of the bar so that the distortion remains stationary in the reference frame of the bar. The growth can be described simply with the mechanism of a driven harmonic oscillator which decreases the vertical frequencies of the stars. At the end of this phase most bar-supporting orbits have banana-like shapes with resonant vertical-to-horizontal frequency ratio close to two. The increase of amplitudes of vertical oscillations leads to the decrease of the amplitudes of horizontal oscillations and the shrinking of the bar. The mass redistribution causes the harmonic oscillators to respond adiabatically and increase the horizontal frequencies. In the following, second phase of buckling, the pattern speed of the distortion increases reaching one third of the circular frequency, but decreases with radius. The distortion propagates as a kinematic bending wave and winds up leaving behind a pronounced boxy/peanut shape. The increased horizontal frequencies cause the weakening of the bar and the transformation of banana-like orbits into pretzel-like ones, except in the outer part of the bar where the banana-like orbits and the distortion survive. The results strongly suggest that the buckling of galactic bars is not related to the fire-hose instability but can be fully explained by the mechanism of vertical resonance creating the distortion that later winds up.
@@ 1 Introduction
Many, perhaps even a majority, of late-type galaxies in the Universe possess an elongated structure in their inner parts known as a bar. For about half a century we have known that galactic disks are generally unstable and form bars spontaneously (Hohl, 1971; Ostriker & Peebles, 1973), and this mechanism has been adopted as the main channel for the formation of bars. Another possibility, identified later on (Gerin et al., 1990; Noguchi, 1996), involves interactions with other galaxies or bigger structures like groups and clusters. Tidal forces naturally present during such encounters lead to the elongation of the stellar component and the formation of the bar if the circumstances of the interactions are favorable in terms of a small enough pericenter distance and prograde configuration (Łokas et al., 2014, 2016; Łokas, 2018; Peschken & Łokas, 2019). Formation and evolution of bars has been studied for decades using $N$-body simulations including later on refinements such as the additional gas component and processes like star formation and feedback. The dark matter halo, originally introduced as a way to moderate the bar instability, has been demonstrated to play a crucial role in bar formation due to its importance in the angular momentum transfer (Athanassoula, 2003).
@@
The simulated bars, after a period of rapid growth, tend to become thicker later on, forming a bulge-like component with a boxy/peanut (BP) shape (Combes & Sanders, 1981; Athanassoula, 2005). This shape has been observed in many spiral galaxies viewed edge-on (Kruk et al., 2019), including our own Milky Way (Weiland et al., 1994; Ciambur et al., 2017). The thickening is often preceded by a short episode of vertical distortion or bending of the bar out of the disk plane, first clearly demonstrated by Raha et al. (1991) and often referred to as buckling instability. In addition to creating the BP shape, buckling also weakens the bar, but generally does not destroy it, and in most configurations it starts to regrow after the event. The instability may occur more than once in a lifetime of a bar, with the second episode usually lasting longer and happening in the outer parts of the bar (Martinez-Valpuesta et al., 2006; Kataria, 2024). It has also been shown in idealized simulations that the presence of the gaseous component tends to suppress buckling (Debattista et al., 2006; Berentzen et al., 2007; Villa-Vargas et al., 2010; Athanassoula, 2013; Łokas, 2020). However, more realistic configurations, modeled in cosmological simulations, do not seem to support this conclusion (Anderson et al., 2024). There is also no obvious correlation between the presence of a BP bulge, probably originating from buckling, and the global gas content in observed galaxies (Erwin et al., 2017).
@@
Figure 1: 2D histograms of the distribution of frequencies along the bar ($f_{x}$) and along the vertical direction ($f_{z}$) for bar-supporting stellar orbits that underwent buckling. The three panels from the left to the right show the distributions before buckling, during buckling and after buckling, respectively. The white dashed lines and formulae indicate the approximate linear relations between the frequencies at the vertical resonance (middle panel) and after buckling (right panel).
Figure 2: Histograms of vertical to horizontal frequency ratios at different stages of the evolution. The blue, green, and red distributions show the values of $f_{z}/f_{x}>2$ (before buckling), $f_{z}^{\prime}/f_{x}\sim 2$ (during buckling) and $f_{z}^{\prime}/f_{x}^{\prime}\leq 2$ (after buckling), respectively.
@@
There are essentially two hypotheses concerning the nature of buckling instability in galactic bars. Both were proposed more than three decades ago and there is still no consensus on which one actually provides a better description of the phenomenon. Combes et al. (1990) and Pfenniger & Friedli (1991) proposed that the instability results from trapping the typical orbits of the bar at vertical resonances. An alternative hypothesis for the origin of buckling involves the ratio of the vertical to horizontal velocity dispersion of the stars in the bar and relates it to the fire-hose instability known from plasma physics (Toomre, 1966; Raha et al., 1991; Merritt & Sellwood, 1994). In this picture, the instability is supposed to be triggered by a sufficiently low ratio of these dispersions, on the order of 0.3, characteristic of strong bars.
@@
In this paper I revisit the issue of the origin and the nature of buckling instability in galactic bars using the insights from recently published works. The next section describes the recent progress that has been made in the field and how it can be used to build a convincing model of buckling instability. Sections 3 and 4 describe two phases of buckling into which the phenomenon can be divided and Section 5 shows how the evolution of the orbital structure of the bar can be understood within this scenario. The discussion follows in Section 6.
Figure 3: Average position of the stars along the vertical direction at times $t=4.1$ to 4.65 Gyr from the start of the simulation. Redder colors indicate the distortion above the disk plane and the bluer ones the distortion below it. The maps show the measurements in the reference frame of the bar, which is oriented along the $x$ axis.
Figure 4: Same as Fig. 3, but for later times, $t=4.7$ to 5.25 Gyr.
@@ 2 Insights from recent studies
In Łokas (2019, hereafter L19) I studied buckling instability using a collisionless simulation of a Milky Way-like galaxy evolving in isolation. The galaxy $N$-body model was initiated with two components: a spherical dark matter halo and an exponential disk, each containing $10^{6}$ particles. The dark matter halo had an Navarro-Frenk-White (Navarro et al., 1997) profile with a virial mass $M_{\rm H}=10^{12}$ M and concentration $c=25$ while the exponential disk had a mass $M_{\rm D}=4.5\times 10^{10}$ M, scale-length $R_{\rm D}=3$ kpc, and thickness $z_{\rm D}=0.42$ kpc. The evolution of the galaxy was followed for 10 Gyr (with outputs saved every 0.05 Gyr) and revealed the bar forming rather slowly, reaching a maximum strength around $t=4.3$ Gyr and the first buckling episode taking place at $t=4.5$ Gyr. The reader is encouraged to browse this earlier paper (L19) because the results presented below refer to and build on these previous findings.
@@
As discussed in L19, in order to study the first buckling event in more detail, the simulation was rerun between $t=3.5$ and $t=5.5$ Gyr saving outputs every 0.001 Gyr (2001 outputs in total) which allows to follow stellar orbits in sufficient detail to measure their properties. The properties of the orbits were estimated in the Cartesian reference frame of the bar with $x$, $y$, and $z$ aligned with the major, intermediate, and minor axes of the bar. Then I calculated the amplitudes of the stellar oscillations along the three axes, $a_{x}$, $a_{y}$, and $a_{z}$, and the frequencies of the orbits, $f_{x}$, $f_{y}$, and $f_{z}$, using discrete Fourier transform in two periods of 0.7 Gyr before and after buckling, that is between 3.5 and 4.2 Gyr and between 4.8 and 5.5 Gyr, respectively. In the following, the quantities measured before buckling will be denoted by symbols without primes while those after buckling will be denoted by the same symbols with primes, so that, for example, $f_{x}$ and $f_{x}^{\prime}$ will correspond to $f_{x,{\rm before}}$ and $f_{x,{\rm after}}$ used in L19. The intermediate period, between 4.2 and 4.8 Gyr, was left out because at these times the orbits vary strongly and the spectral analysis of stellar orbits is not reliable. I emphasize that the measurements of orbital properties were done “in vivo”, that is in the live evolving bar rather than in a frozen potential, as is often done in the literature. Selecting bar-supporting orbits with $a_{x}<7$ kpc (smaller than the length of the bar), $a_{y}<0.7a_{x}$ (elongated along the bar) and those that underwent buckling, that is those that increased their $a_{z}$ amplitude during the evolution, I ended up with a sample of almost $2\times 10^{5}$ orbits that were studied further.
Figure 5: Getting to the resonance. The three columns show three examples of the evolution of stellar orbits under the force due to the existing distortion for different combinations of the initial frequency ($\omega$) and the force amplitude $F_{0}$: $(\omega,F_{0})=$ (2.5, 3), (3, 6) and (3.5, 10). The blue lines correspond to the initial oscillations and orbits and the red lines show the modified ones. The panels of the upper row show the variation of the vertical position of the star along the orbit while the next rows show the full orbits combining the motion along $z$ with an ellipse in the $xy$ plane before (blue, second row), after (red, third row) the action of the force and the two combined (lower row).
@@
The most important finding of that study was the discovery of a very tight linear relation between the orbital frequencies of the stellar orbits after buckling
$f_{z}^{\prime}=(4/3)(f_{x}^{\prime}+f_{\rm p}),$ (1)
@@
where $f_{\rm p}$ was identified numerically as the pattern speed of the bar, which did not vary strongly during buckling and for this particular simulation was equal to $f_{\rm p}=2.5$ Gyr-1. It might be useful to note that in the more commonly used notation, with the circular frequency $\Omega^{\prime}=f_{x}^{\prime}+f_{\rm p}$ and the vertical frequency $\nu^{\prime}=f_{z}^{\prime}$ this relation translates to an even simpler formula: $3\nu^{\prime}=4\Omega^{\prime}$. The relation, in terms of the 2D histogram of the distribution of the frequencies and the line given by equation (1) are shown in the right panel of Fig. 1.
@@
The existence of this very tight relation between the vertical and horizontal frequencies after buckling was later confirmed by Sellwood & Gerhard (2020), Li et al. (2023), and most recently by McClure et al. (2024) (for the case without the initial bulge), although they did not acknowledge the role of the pattern speed of the bar in it. However, using their published results concerning the appearance of the relation and the pattern speed of their simulated bars I confirmed that the relation (1) is also obeyed by their results. It is especially noteworthy given that the pattern speeds of the bars are different in all these simulations. This confirms that the presence of the pattern speed in this relation is not just a numerical coincidence but has a deeper meaning.
@@
Comparing this relation with the much wider distribution of frequencies before buckling (left panel of Fig. 1) one can see that after buckling the relation becomes much tighter and its slope changes. While before buckling the distribution barely touches the $f_{z}=2f_{x}$ region, after buckling all the stars are below it. This strongly suggests that they must have crossed the vertical 2:1 resonance, that is the $f_{z}=2f_{x}$ line at a certain time. It is difficult to verify this using the spectral analysis of stellar orbits because the latter has to be performed on time-spans containing some number of oscillations. Instead, Li et al. (2023) have recently managed to determine the instantaneous frequencies of stellar orbits in their simulation of a buckling bar by measuring directly their periods. They confirmed that during buckling majority of the stellar orbits in the bar indeed cross the 2:1 vertical resonance and thus provided strong arguments in favor of this mechanism as the one responsible for triggering the vertical distortion.
@@
Figure 6: Average orbit of stars at the time of the vertical resonance, in the first stage of buckling, viewed along three directions: edge-on (upper panel), intermediate (middle panel) and face-on (lower panel). The dashed line indicates the more distant part of the orbit in the edge-on view.
@@
Figure 7: Mean velocity of the stars along $x$ (upper panel) and along $z$ (lower panel) in the bar at the time of maximum distortion, $t=4.5$ Gyr, at the end of the first phase of buckling. The white arrows indicate the bulk motion of the stars in a given direction and the small black arrows on the orbit (black line) indicate the direction of motion of the star along the orbit.
@@
Interestingly, plotting $f_{z}^{\prime}$ (after buckling) as a function of $f_{x}$ (before buckling) we still get approximately the 2:1 resonance (middle panel of Fig. 1). Does it mean that on the way to the resonance only $f_{z}$ changes while $f_{x}$ remains approximately constant? Yes, because if the resonance is present even if $f_{x}$ is not yet changed then it means that only the variation of $f_{z}$ is responsible for getting to the resonance! It does not mean that $f_{x}$ remains constant during the whole evolution, but only that it changes mostly after the resonance is reached. This observation suggests a simple scenario, in which the buckling event can be divided into two distinct phases: the first one in which $f_{z}$ decreases while $f_{x}$ remains approximately constant and the second one in which $f_{x}$ increases while $f_{z}$ remains unchanged. The next two sections describe the two phases in more detail.
@@
Before moving on it may be interesting to look at the evolution of the frequencies once again, but this time in the form of histograms of the ratio $f_{z}/f_{x}$, often discussed in similar studies (Portail et al., 2015). Figure 2 shows how the ratio evolved, from $f_{z}/f_{x}>2$ before buckling (blue), to $f_{z}^{\prime}/f_{x}\sim 2$ during buckling (green), and $f_{z}^{\prime}/f_{x}^{\prime}\leq 2$ after buckling (red). Let me recall that the green histogram shows a combination of the frequencies measured before buckling ($f_{x}$) and after buckling ($f_{z}^{\prime}$) rather than real frequency estimates during buckling, but still the distribution, although rather wide, peaks very close to the value of two expected for the vertical resonance.

### v2

@@
Many strong simulated galactic bars experience buckling instability, which manifests itself as a vertical distortion out of the disk plane, and later dissipates. Using a simulation of an isolated Milky Way-like galaxy, I demonstrate that the phenomenon can be divided into two distinct phases. In the first one, the distortion grows and its pattern speed remains equal to the pattern speed of the bar, so that the distortion remains stationary in the reference frame of the bar. The growth can be described with the mechanism of a driven harmonic oscillator with time-dependent force, which decreases the vertical frequencies of the stars. At the end of this phase, most bar-supporting orbits have banana-like shapes with a resonant vertical-to-horizontal frequency ratio close to two. The increase of amplitudes of vertical oscillations leads to the decrease of the amplitudes of horizontal oscillations and the shrinking of the bar. The mass redistribution causes the harmonic oscillators to respond adiabatically and increase the horizontal frequencies. In the following second phase of buckling, the pattern speed of the distortion increases – reaching one third of the circular frequency – but it decreases with radius. The distortion propagates as a kinematic bending wave and winds up, leaving behind a pronounced boxy/peanut shape. The increased horizontal frequencies cause the weakening of the bar and the transformation of banana-like orbits into pretzel-like ones, except in the outer part of the bar, where the banana-like orbits and the distortion survive. The results strongly suggest that the buckling of galactic bars is not related to the fire-hose instability, but it can be fully explained by the mechanism of vertical resonance creating the distortion that later winds up.
@@ 1 Introduction
Many, perhaps even a majority, of late-type galaxies in the Universe possess an elongated structure in their inner parts, which is known as a bar. For about half a century we have known that galactic disks are generally unstable and form bars spontaneously (Hohl, 1971; Ostriker & Peebles, 1973), and this mechanism has been adopted as the main channel for the formation of bars. Another possibility, identified later on (Gerin et al., 1990; Noguchi, 1996), involves interactions with other galaxies or bigger structures such as groups and clusters. Tidal forces naturally present during such encounters lead to the elongation of the stellar component and the formation of the bar if the circumstances of the interactions are favorable in terms of a small enough pericenter distance and prograde configuration (Łokas et al., 2014, 2016; Łokas, 2018; Peschken & Łokas, 2019). The formation and evolution of bars has been studied for decades using $N$-body simulations, and later on refinements such as the additional gas component and processes such as star formation and feedback. The dark-matter halo, originally introduced as a way to moderate the bar instability, has been shown to play a crucial role in bar formation due to its importance in the transfer of angular momentum (Athanassoula, 2003).
@@
The simulated bars, after a period of rapid growth, tend to become thicker later on, forming a bulge-like component with a boxy/peanut (BP) shape (Combes & Sanders, 1981; Athanassoula, 2005). This shape has been observed in many spiral galaxies viewed edge-on (Kruk et al., 2019), including our own Milky Way (Weiland et al., 1994; Ciambur et al., 2017). The thickening is often preceded by a short episode of vertical distortion or bending of the bar out of the disk plane; it was first described by Friedli & Pfenniger (1990), Pfenniger & Friedli (1991), and Raha et al. (1991) and is often referred to as buckling instability. However, it should be noted that the BP shape can also form without buckling (Combes et al., 1990; Pfenniger & Friedli, 1991; Sellwood & Gerhard, 2020; Li et al., 2023).
In addition to creating the BP shape, buckling also weakens the bar, but generally does not destroy it, and in most configurations it starts to regrow after the event. The instability may occur more than once in the lifetime of a bar, with the second episode usually lasting longer and happening in the outer parts (Martinez-Valpuesta et al., 2006; Kataria, 2024). It has also been shown in idealized simulations that the presence of the gaseous component tends to suppress buckling (Debattista et al., 2006; Berentzen et al., 2007; Villa-Vargas et al., 2010; Athanassoula, 2013; Łokas, 2020). However, more realistic configurations, modeled in cosmological simulations, do not seem to support this conclusion (Anderson et al., 2024). There is also no obvious correlation between the presence of a BP bulge, probably originating from buckling, and the global gas content in observed galaxies (Erwin et al., 2017).
@@
Figure 1: 2D histograms of distribution of frequencies along the bar ($f_{x}$) and along the vertical direction ($f_{z}$) for bar-supporting stellar orbits that underwent buckling. The three panels from the left to the right show the distributions before buckling, during buckling, and after buckling, respectively. The white dashed lines and formulae indicate the approximate linear relations between the frequencies at the vertical resonance (middle panel) and after buckling (right panel).
@@
Two hypotheses were proposed concerning the nature of buckling instability in galactic bars. Combes et al. (1990) and Pfenniger & Friedli (1991) suggested that the formation of the BP shape results from the trapping of typical orbits of the bar at vertical resonances. An alternative hypothesis for the origin of buckling involves the ratio of the vertical-to-horizontal velocity dispersion of the stars in the bar and relates it to the fire-hose instability known from plasma physics (Toomre, 1966; Raha et al., 1991; Merritt & Sellwood, 1994). In this picture, the instability is supposed to be triggered by a sufficiently low ratio of these dispersions (on the order of 0.3), which is characteristic of strong bars. Recently, the hypothesis based on vertical resonance seems to be getting more support from numerical studies (Li et al., 2023; McClure et al., 2025).
@@
In this work, I revisited the issue of the origin and the nature of buckling instability in galactic bars using the insights from recently published works. The next section describes the recent progress that has been made in the field and how it can be used to build a convincing model of buckling instability. Sections 3 and 4 describe two phases of buckling into which the phenomenon can be divided, and Section 5 shows how the evolution of the orbital structure of the bar can be understood within this scenario. The discussion follows in Section 6.
@@ 2 Insights from recent studies
In Łokas (2019, hereafter L19), I studied buckling instability using a collisionless simulation of a Milky Way-like galaxy evolving in isolation. The galaxy $N$-body model was initiated with two components: a spherical dark-matter halo and an exponential disk, each containing $10^{6}$ particles. The dark-matter halo had a Navarro-Frenk-White (Navarro et al., 1997) profile with a virial mass of $M_{\rm H}=10^{12}$ M and concentration of $c=25,$ while the exponential disk had a mass of $M_{\rm D}=4.5\times 10^{10}$ M, scale-length of $R_{\rm D}=3$ kpc, and thickness of $z_{\rm D}=0.42$ kpc. The evolution of the galaxy was followed for 10 Gyr (with outputs saved every 0.05 Gyr) and revealed the bar to be forming rather slowly, reaching a maximum strength around $t=4.3$ Gyr and the first buckling episode taking place at $t=4.5$ Gyr. The reader is encouraged to browse this earlier paper (L19), because the results presented below refer to and build on previous findings.
@@
As discussed in L19, in order to study the first buckling event in more detail, the simulation was rerun between $t=3.5$ and $t=5.5$ Gyr, saving outputs every 0.001 Gyr (2001 outputs in total), which allowed me to follow stellar orbits in sufficient detail to measure their properties. The properties of the orbits were estimated in the Cartesian reference frame of the bar with $x$, $y$, and $z$ aligned with the major, intermediate, and minor axes of the bar. Then, I calculated the amplitudes of the stellar oscillations along the three axes, $a_{x}$, $a_{y}$, and $a_{z}$, and the frequencies of the orbits, $f_{x}$, $f_{y}$, and $f_{z}$, using a discrete Fourier transform in two periods of 0.7 Gyr before and after buckling; that is, between 3.5 and 4.2 Gyr and between 4.8 and 5.5 Gyr, respectively. In the following, the quantities measured before buckling are denoted by symbols without primes, while those measured after buckling are denoted by the same symbols with primes; hence, $f_{x}$ and $f_{x}^{\prime}$ will correspond to $f_{x,{\rm before}}$ and $f_{x,{\rm after}}$ used in L19, for example. The intermediate period, between 4.2 and 4.8 Gyr, was left out because at these times the orbits vary strongly, and the spectral analysis of stellar orbits is not reliable. It should be emphasized that the measurements of orbital properties were done “in vivo”, that is, in the live evolving bar rather than in a frozen potential, as is often done in the literature. Selecting bar-supporting orbits with $a_{x}<7$ kpc (smaller than the length of the bar), $a_{y}<0.7a_{x}$ (elongated along the bar), and those that underwent buckling – that is, those that increased their $a_{z}$ amplitude during the evolution – gave a sample of almost $2\times 10^{5}$ orbits that were studied further.
@@
The most important finding of that study was the discovery of a very tight linear relation between the orbital frequencies of the stellar orbits after buckling, which is very well fit (with the correlation coefficient of 0.99) by the relation $f_{z}^{\prime}=1.35f_{x}^{\prime}+3.39$ and can be approximated by the formula
$f_{z}^{\prime}=(4/3)(f_{x}^{\prime}+f_{\rm p}).$ (1)
@@
The parameter $f_{\rm p}$ was identified numerically as the pattern speed of the bar, which did not vary strongly during buckling, and for this particular simulation it was equal to $f_{\rm p}=2.5$ Gyr-1. It might be useful to note that in the more commonly used notation, with the circular frequency $\Omega^{\prime}=f_{x}^{\prime}+f_{\rm p}$ and the vertical frequency $\nu^{\prime}=f_{z}^{\prime}$, this relation translates to an even simpler formula: $3\nu^{\prime}=4\Omega^{\prime}$. The relation, in terms of the 2D histogram of the distribution of the frequencies and the line given by Equation (1), is shown in the right panel of Fig. 1. Its origin is explored further below and in Section 4, where its connection to the pattern speed of the vertical distortion is described.
@@
The existence of this very tight relation between the vertical and horizontal frequencies after buckling was later confirmed by Sellwood & Gerhard (2020), Li et al. (2023), and most recently by McClure et al. (2025) (for the case without the initial bulge), although they did not acknowledge the role of the pattern speed of the bar. However, using their published results concerning the appearance of the relation and the pattern speed of their simulated bars, one can easily demonstrate that the relation (1) is also obeyed by their results. It is especially noteworthy given that the pattern speeds of the bars are different in all these simulations. This confirms that the presence of the pattern speed in this relation is not just a numerical coincidence, it has a deeper meaning.
@@
Comparing this relation with the much wider distribution of frequencies before buckling (left panel of Fig. 1), one can see that after buckling the relation becomes much tighter, and its slope changes significantly. While before buckling the distribution barely touches the $f_{z}=2f_{x}$ region (from above), after buckling all the stars are well below it. This strongly suggests that they must have crossed the vertical 2:1 resonance – that is, the $f_{z}=2f_{x}$ line – at a certain time. It is difficult to verify this using the spectral analysis of stellar orbits because the latter has to be performed on time spans containing many oscillations. Instead, Li et al. (2023) recently managed to determine the instantaneous frequencies of stellar orbits in their simulation of a buckling bar by measuring their periods directly. They confirmed that during buckling, the majority of stellar orbits in the bar indeed cross the 2:1 vertical resonance and thus provide strong arguments in favor of this mechanism as that responsible for triggering the vertical distortion.
Figure 2: Histograms of vertical-to-horizontal frequency ratios at different stages of evolution. The blue, green, and red distributions show the values of $f_{z}/f_{x}>2$ (before buckling), $f_{z}^{\prime}/f_{x}\sim 2$ (during buckling), and $f_{z}^{\prime}/f_{x}^{\prime}\leq 2$ (after buckling), respectively.
Figure 3: Average position of stars along vertical direction at times $t=4.1$ to 4.65 Gyr from the start of the simulation. Redder colors indicate the distortion above the disk plane, and the bluer ones show the distortion below it. The maps show the measurements in the reference frame of the bar, which is oriented along the $x$-axis.
Figure 4: Reaching resonance. The three columns show three examples of the evolution of stellar orbits under the growing force due to the distortion for different values of the initial frequency: $\omega=2.5$, 3, and 3.5. The solutions to Equation (2) were obtained with the force parameters $F_{0}=10$ and $\alpha=0.1$ in all cases. The blue lines correspond to the initial oscillations and orbits, while the red lines show the modified ones. The panels of the upper row show the variation of the vertical position of the star along the orbit, while the next rows show the full orbits combining the motion along $z$ with an ellipse in the $xy$ plane without (blue, second row) and with (red, third row) the action of the force and the two combined (lower row).
@@
Figure 5: Average orbit of stars at time of vertical resonance, in first stage of buckling, viewed along three directions: edge-on (upper panel), intermediate (middle panel), and face-on (lower panel). The dashed line indicates the more distant part of the orbit in the edge-on view.
@@
Figure 6: Mean velocity of stars along $x$ (upper panel) and along $z$ (lower panel) in the bar at the time of maximum distortion, $t=4.5$ Gyr, at the end of the first phase of buckling. The white arrows indicate the bulk motion of the stars in a given direction, and the small black arrows on the orbit (black line) indicate the direction of motion of the star along the orbit.
@@
Interestingly, plotting $f_{z}^{\prime}$ (after buckling) as a function of $f_{x}$ (before buckling), one still obtains approximately the 2:1 resonance (middle panel of Fig. 1). This means that on the way to the resonance only $f_{z}$ changes, while $f_{x}$ remains approximately constant; if the resonance is present, even if $f_{x}$ is not yet changed it means that only the variation of $f_{z}$ is responsible for getting to the resonance. It does not mean that $f_{x}$ remains constant during the whole evolution, only that it changes mostly after the resonance is reached. This observation suggests a simple scenario, in which the buckling event can be divided into two distinct phases: the first one, where $f_{z}$ decreases while $f_{x}$ remains approximately constant; and the second one, where $f_{x}$ increases while $f_{z}$ remains unchanged. The next two sections describe the two phases in more detail.
@@
Before moving on, it may be interesting to look at the evolution of the frequencies once again, but this time in the form of histograms of the ratio $f_{z}/f_{x}$, often discussed in similar studies (Portail et al., 2015). Figure 2 shows how the ratio evolved, from $f_{z}/f_{x}>2$ before buckling (blue), to $f_{z}^{\prime}/f_{x}\sim 2$ during buckling (green), and $f_{z}^{\prime}/f_{x}^{\prime}\leq 2$ after buckling (red). I remind the reader that the green histogram shows a combination of the frequencies measured before buckling ($f_{x}$) and after buckling ($f_{z}^{\prime}$), rather than real frequency estimates during buckling; however, the distribution, although rather wide, still peaks very close to the value of two expected for the vertical resonance.